# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 9. Acknowledgments
- Inan et al. (2023) Hakan Inan, Kartikeya Upasani, Jianfeng Chi, Rashi Rungta, Krithika Iyer, Yuning Mao, Michael Tontchev, Qing Hu, Brian Fuller, Davide Testuggine, et al. 2023. Llama guard: Llm-based input-output safeguard for human-ai conversations. arXiv preprint arXiv:2312.06674 (2023).
- Ji et al. (2024a) Jiaming Ji, Mickel Liu, Josef Dai, Xuehai Pan, Chi Zhang, Ce Bian, Boyuan Chen, Ruiyang Sun, Yizhou Wang, and Yaodong Yang. 2024a. Beavertails: Towards improved safety alignment of llm via a human-preference dataset. NIPS (2024).
- Ji et al. (2024b) Yuheng Ji, Yue Liu, Zhicheng Zhang, Zhao Zhang, Yuting Zhao, Gang Zhou, Xingwei Zhang, Xinwang Liu, and Xiaolong Zheng. 2024b. AdvLoRA: Adversarial Low-Rank Adaptation of Vision-Language Models. arXiv preprint arXiv:2404.13425 (2024).
- Jia et al. (2019) Xiaojun Jia, Xingxing Wei, Xiaochun Cao, and Hassan Foroosh. 2019. Comdefend: An efficient image compression model to defend adversarial examples. In CVPR.
- Jia et al. (2020) Xiaojun Jia, Xingxing Wei, Xiaochun Cao, and Xiaoguang Han. 2020. Adv-watermark: A novel watermark perturbation for adversarial examples. In ACM MM.
- Jiang et al. (2023) Albert Q Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Florian Bressand, Gianna Lengyel, Guillaume Lample, Lucile Saulnier, et al. 2023. Mistral 7B. arXiv preprint arXiv:2310.06825 (2023).
- Jigsaw (2017) Google Jigsaw. 2017. Perspective API. (2017). https://www.perspectiveapi.com/
- Jin et al. (2024) Haibo Jin, Leyang Hu, Xinuo Li, Peiyan Zhang, Chonghan Chen, Jun Zhuang, and Haohan Wang. 2024. Jailbreakzoo: Survey, landscapes, and horizons in jailbreaking large language and vision-language models. arXiv preprint arXiv:2407.01599 (2024).
- Jin et al. (2021) Woojeong Jin, Yu Cheng, Yelong Shen, Weizhu Chen, and Xiang Ren. 2021. A good prompt is worth millions of parameters: Low-resource prompt-based learning for vision-language models. arXiv preprint arXiv:2110.08484 (2021).
@@
- Kannan et al. (2018) Harini Kannan, Alexey Kurakin, and Ian Goodfellow. 2018. Adversarial logit pairing. arXiv preprint arXiv:1803.06373 (2018).
- Kazemzadeh et al. (2014) Sahar Kazemzadeh, Vicente Ordonez, Mark Matten, and Tamara Berg. 2014. Referitgame: Referring to objects in photographs of natural scenes. In EMNLP.
@@
- Kingma (2013) Diederik P Kingma. 2013. Auto-encoding variational bayes. arXiv preprint arXiv:1312.6114 (2013).
- Kingma (2014) Diederik P Kingma. 2014. Adam: A method for stochastic optimization. arXiv preprint arXiv:1412.6980 (2014).
@@
- Krizhevsky et al. (2009) Alex Krizhevsky, Geoffrey Hinton, et al. 2009. Learning multiple layers of features from tiny images. (2009).
@@
- Kurakin et al. (2016) Alexey Kurakin, Ian Goodfellow, and Samy Bengio. 2016. Adversarial machine learning at scale. arXiv preprint arXiv:1611.01236 (2016).
- Kurakin et al. (2018) Alexey Kurakin, Ian J Goodfellow, and Samy Bengio. 2018. Adversarial examples in the physical world. In Artificial intelligence safety and security. Chapman and Hall/CRC.
@@
- Laidlaw et al. (2020) Cassidy Laidlaw, Sahil Singla, and Soheil Feizi. 2020. Perceptual adversarial robustness: Defense against unseen threat models. arXiv preprint arXiv:2006.12655 (2020).
- LAION (2023) LAION. 2023. Reaching 80 zero-shot accuracy with openclip: Vit-g/14 trained on laion-2b. (2023). https://laion.ai/blog/giant-openclip/
- Lan et al. (2019) Xiangyuan Lan, Mang Ye, Rui Shao, Bineng Zhong, Pong C Yuen, and Huiyu Zhou. 2019. Learning modality-consistency feature templates: A robust RGB-infrared tracking system. IEEE Transactions on Industrial Electronics (2019).
- Larsen et al. (2016) Anders Boesen Lindbo Larsen, Søren Kaae Sønderby, Hugo Larochelle, and Ole Winther. 2016. Autoencoding beyond pixels using a learned similarity metric. In ICML.
@@
- LeCun et al. (1998) Yann LeCun, Léon Bottou, Yoshua Bengio, and Patrick Haffner. 1998. Gradient-based learning applied to document recognition. Proc. IEEE (1998).
- Lees et al. (2022) Alyssa Lees, Vinh Q Tran, Yi Tay, Jeffrey Sorensen, Jai Gupta, Donald Metzler, and Lucy Vasserman. 2022. A new generation of perspective api: Efficient multilingual character-level transformers. In SIGKDD.
- Li et al. (2023b) Bohao Li, Rui Wang, Guangzhi Wang, Yuying Ge, Yixiao Ge, and Ying Shan. 2023b. Seed-bench: Benchmarking multimodal llms with generative comprehension. arXiv preprint arXiv:2307.16125 (2023).
- Li et al. (2023c) Bo Li, Yuanhan Zhang, Liangyu Chen, Jinghao Wang, Jingkang Yang, and Ziwei Liu. 2023c. Otter: A Multi-Modal Model with In-Context Instruction Tuning. CoRR abs/2305.03726 (2023).
- Li et al. (2018) Hao Li, Zheng Xu, Gavin Taylor, Christoph Studer, and Tom Goldstein. 2018. Visualizing the loss landscape of neural nets. NIPS (2018).
- Li et al. (2023a) Junnan Li, Dongxu Li, Silvio Savarese, and Steven Hoi. 2023a. Blip-2: Bootstrapping language-image pre-training with frozen image encoders and large language models. In ICML.
- Li et al. (2022) Junnan Li, Dongxu Li, Caiming Xiong, and Steven Hoi. 2022. Blip: Bootstrapping language-image pre-training for unified vision-language understanding and generation. In ICML.
- Li et al. (2019) Juncheng Li, Frank Schmidt, and Zico Kolter. 2019. Adversarial camera stickers: A physical camera-based attack on deep learning systems. In ICML.
- Li et al. (2021) Junnan Li, Ramprasaath Selvaraju, Akhilesh Gotmare, Shafiq Joty, Caiming Xiong, and Steven Chu Hong Hoi. 2021. Align before fuse: Vision and language representation learning with momentum distillation. NIPS (2021).
- Li et al. (2024a) Lijun Li, Bowen Dong, Ruohui Wang, Xuhao Hu, Wangmeng Zuo, Dahua Lin, Yu Qiao, and Jing Shao. 2024a. Salad-bench: A hierarchical and comprehensive safety benchmark for large language models. arXiv preprint arXiv:2402.05044 (2024).
- Li et al. (2024b) Lin Li, Haoyan Guan, Jianing Qiu, and Michael Spratling. 2024b. One prompt word is enough to boost adversarial robustness for pre-trained vision-language models. In CVPR.
- Li et al. (2020b) Maosen Li, Cheng Deng, Tengjiao Li, Junchi Yan, Xinbo Gao, and Heng Huang. 2020b. Towards transferable targeted attack. In CVPR.
- Li et al. (2024d) Mukai Li, Lei Li, Yuwei Yin, Masood Ahmed, Zhenguang Liu, and Qi Liu. 2024d. Red teaming visual language models. arXiv preprint arXiv:2401.12915 (2024).
@@
- Li et al. (2020a) Yingwei Li, Song Bai, Cihang Xie, Zhenyu Liao, Xiaohui Shen, and Alan Yuille. 2020a. Regional homogeneity: Towards learning transferable universal adversarial perturbations against defenses. In ECCV.
- Li et al. (2024c) Yifan Li, Hangyu Guo, Kun Zhou, Wayne Xin Zhao, and Ji-Rong Wen. 2024c. Images are Achilles’ Heel of Alignment: Exploiting Visual Vulnerabilities for Jailbreaking Multimodal Large Language Models. arXiv preprint arXiv:2403.09792 (2024).
- Liao et al. (2018) Fangzhou Liao, Ming Liang, Yinpeng Dong, Tianyu Pang, Xiaolin Hu, and Jun Zhu. 2018. Defense against adversarial attacks using high-level representation guided denoiser. In CVPR.
@@
- Lin et al. (2019) Jiadong Lin, Chuanbiao Song, Kun He, Liwei Wang, and John E Hopcroft. 2019. Nesterov accelerated gradient and scale invariance for adversarial attacks. arXiv preprint arXiv:1908.06281 (2019).
- Lin et al. (2017) Yen-Chen Lin, Zhang-Wei Hong, Yuan-Hong Liao, Meng-Li Shih, Ming-Yu Liu, and Min Sun. 2017. Tactics of adversarial attack on deep reinforcement learning agents. arXiv preprint arXiv:1703.06748 (2017).
- Lin et al. (2023) Zi Lin, Zihan Wang, Yongqi Tong, Yangkun Wang, Yuxin Guo, Yujia Wang, and Jingbo Shang. 2023. Toxicchat: Unveiling hidden challenges of toxicity detection in real-world user-ai conversation. arXiv preprint arXiv:2310.17389 (2023).
- Liu et al. (2020) Aishan Liu, Jiakai Wang, Xianglong Liu, Bowen Cao, Chongzhi Zhang, and Hang Yu. 2020. Bias-based universal adversarial patch attack for automatic check-out. In ECCV.
- Liu et al. (2024d) Daizong Liu, Mingyu Yang, Xiaoye Qu, Pan Zhou, Wei Hu, and Yu Cheng. 2024d. A survey of attacks on large vision-language models: Resources, advances, and future trends. arXiv preprint arXiv:2407.07403 (2024).
- Liu et al. (2023c) Fangcheng Liu, Chao Zhang, and Hongyang Zhang. 2023c. Towards transferable unrestricted adversarial examples with minimum changes. In SaTML.
- Liu et al. (2024a) Haotian Liu, Chunyuan Li, Yuheng Li, and Yong Jae Lee. 2024a. Improved baselines with visual instruction tuning. In CVPR.
- Liu et al. (2024b) Haotian Liu, Chunyuan Li, Yuheng Li, Bo Li, Yuanhan Zhang, Sheng Shen, and Yong Jae Lee. 2024b. Llava-next: Improved reasoning, ocr, and world knowledge.
- Liu et al. (2024c) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2024c. Visual instruction tuning. NIPS (2024).
- Liu et al. (2023d) X Liu, Y Zhu, J Gu, Y Lan, C Yang, and Y Qiao. 2023d. Mm-safetybench: A benchmark for safety evaluation of multimodal large language models. arXiv preprint arXiv:2311.17600 (2023).
- Liu (2019) Yinhan Liu. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692 (2019).
- Liu et al. (2016) Yanpei Liu, Xinyun Chen, Chang Liu, and Dawn Song. 2016. Delving into transferable adversarial examples and black-box attacks. arXiv preprint arXiv:1611.02770 (2016).
- Liu et al. (2022) Ye Liu, Yaya Cheng, Lianli Gao, Xianglong Liu, Qilong Zhang, and Jingkuan Song. 2022. Practical evaluation of adversarial robustness via adaptive auto attack. In CVPR.
- Liu et al. (2023a) Yi Liu, Gelei Deng, Zhengzi Xu, Yuekang Li, Yaowen Zheng, Ying Zhang, Lida Zhao, Tianwei Zhang, Kailong Wang, and Yang Liu. 2023a. Jailbreaking chatgpt via prompt engineering: An empirical study. arXiv preprint arXiv:2305.13860 (2023).
- Liu et al. (2023b) Yuan Liu, Haodong Duan, Yuanhan Zhang, Bo Li, Songyang Zhang, Wangbo Zhao, Yike Yuan, Jiaqi Wang, Conghui He, Ziwei Liu, et al. 2023b. Mmbench: Is your multi-modal model an all-around player? arXiv preprint arXiv:2307.06281 (2023).
- Liu et al. (2021) Ze Liu, Yutong Lin, Yue Cao, Han Hu, Yixuan Wei, Zheng Zhang, Stephen Lin, and Baining Guo. 2021. Swin transformer: Hierarchical vision transformer using shifted windows. In ICCV.
- Liu et al. (2019) Zihao Liu, Qi Liu, Tao Liu, Nuo Xu, Xue Lin, Yanzhi Wang, and Wujie Wen. 2019. Feature distillation: Dnn-oriented jpeg compression against adversarial examples. In CVPR.
@@
- LMSYS (2023) LMSYS. 2023. Vicuna-7b-v1.5. (2023). https://huggingface.co/lmsys/vicuna-7b-v1.5
- Long et al. (2022) Yuyang Long, Qilong Zhang, Boheng Zeng, Lianli Gao, Xianglong Liu, Jian Zhang, and Jingkuan Song. 2022. Frequency domain model augmentation for adversarial attack. In ECCV.
- Lu et al. (2024) Dong Lu, Tianyu Pang, Chao Du, Qian Liu, Xianjun Yang, and Min Lin. 2024. Test-time backdoor attacks on multimodal large language models. arXiv preprint arXiv:2402.08577 (2024).
- Luo et al. (2024a) Haochen Luo, Jindong Gu, Fengyuan Liu, and Philip Torr. 2024a. An image is worth 1000 lies: Adversarial transferability across prompts on vision-language models. arXiv preprint arXiv:2403.09766 (2024).
- Luo et al. (2024b) Weidi Luo, Siyuan Ma, Xiaogeng Liu, Xiaoyu Guo, and Chaowei Xiao. 2024b. Jailbreakv-28k: A benchmark for assessing the robustness of multimodal large language models against jailbreak attacks. arXiv preprint arXiv:2404.03027 (2024).
- Ma et al. (2024) Siyuan Ma, Weidi Luo, Yu Wang, Xiaogeng Liu, Muhao Chen, Bo Li, and Chaowei Xiao. 2024. Visual-RolePlay: Universal Jailbreak Attack on MultiModal Large Language Models via Role-playing Image Characte. arXiv preprint arXiv:2405.20773 (2024).
- Madry (2017) Aleksander Madry. 2017. Towards deep learning models resistant to adversarial attacks. arXiv preprint arXiv:1706.06083 (2017).
- Mao et al. (2016) Junhua Mao, Jonathan Huang, Alexander Toshev, Oana Camburu, Alan L Yuille, and Kevin Murphy. 2016. Generation and comprehension of unambiguous object descriptions. In CVPR.
@@
- Marino et al. (2019) Kenneth Marino, Mohammad Rastegari, Ali Farhadi, and Roozbeh Mottaghi. 2019. Ok-vqa: A visual question answering benchmark requiring external knowledge. In CVPR.
- Mazeika et al. (2024) Mantas Mazeika, Long Phan, Xuwang Yin, Andy Zou, Zifan Wang, Norman Mu, Elham Sakhaee, Nathaniel Li, Steven Basart, Bo Li, et al. 2024. Harmbench: A standardized evaluation framework for automated red teaming and robust refusal. arXiv preprint arXiv:2402.04249 (2024).
- Mazeika and Forsyth (2023) Zou A. Mu N. Phan L. Wang Z. Yu C. Khoja A. Jiang F. O’Gara A. Sakhaee E. Xiang Z. Rajabi A. Hendrycks D. Poovendran R. Li B. Mazeika, M. and D Forsyth. 2023. Tdc 2023 (llm edition): The trojan detection challenge. In NeurIPS Competition Track.
- Microsoft (2023a) Microsoft. 2023a. Bing chat. (2023). https://www.bing.com/new
- Microsoft (2023b) Microsoft. 2023b. Copilot. (2023). https://copilot.microsoft.com/
- Mogelmose et al. (2012) Andreas Mogelmose, Mohan Manubhai Trivedi, and Thomas B Moeslund. 2012. Vision-based traffic sign detection and analysis for intelligent driver assistance systems: Perspectives and survey. TITS (2012).
@@
- Moosavi-Dezfooli et al. (2016) Seyed-Mohsen Moosavi-Dezfooli, Alhussein Fawzi, and Pascal Frossard. 2016. Deepfool: a simple and accurate method to fool deep neural networks. In CVPR.
@@
- Narodytska and Kasiviswanathan (2016) Nina Narodytska and Shiva Prasad Kasiviswanathan. 2016. Simple black-box adversarial perturbations for deep networks. arXiv preprint arXiv:1612.06299 (2016).
@@
- Naseer et al. (2020) Muzammal Naseer, Salman Khan, Munawar Hayat, Fahad Shahbaz Khan, and Fatih Porikli. 2020. A self-supervised approach for adversarial robustness. In CVPR.
@@
- Nasr et al. (2023) Milad Nasr, Nicholas Carlini, Jonathan Hayase, Matthew Jagielski, A Feder Cooper, Daphne Ippolito, Christopher A Choquette-Choo, Eric Wallace, Florian Tramèr, and Katherine Lee. 2023. Scalable extraction of training data from (production) language models. arXiv preprint arXiv:2311.17035 (2023).
@@
- Netzer et al. (2011) Yuval Netzer, Tao Wang, Adam Coates, Alessandro Bissacco, Baolin Wu, Andrew Y Ng, et al. 2011. Reading digits in natural images with unsupervised feature learning. In NIPS Workshop.
@@
- Nguyen et al. (2015) Anh Nguyen, Jason Yosinski, and Jeff Clune. 2015. Deep neural networks are easily fooled: High confidence predictions for unrecognizable images. In CVPR.
@@
- Nie et al. (2022) Weili Nie, Brandon Guo, Yujia Huang, Chaowei Xiao, Arash Vahdat, and Anima Anandkumar. 2022. Diffusion models for adversarial purification. arXiv preprint arXiv:2205.07460 (2022).
- NIPS (2017) NIPS. 2017. NIPS 2017 Competition Track. (2017). https://nips.cc/Conferences/2017/CompetitionTrack

### v2

@@ 9. Acknowledgments
- Inan et al. (2023) Hakan Inan, Kartikeya Upasani, et al. 2023. Llama guard: Llm-based input-output safeguard for human-ai conversations. arXiv:2312.06674.
- Ji et al. (2024a) Jiaming Ji, Mickel Liu, Josef Dai, et al. 2024a. Beavertails: Towards improved safety alignment of llm via a human-preference dataset. In NIPS.
- Ji et al. (2024b) Yuheng Ji, Yue Liu, et al. 2024b. AdvLoRA: Adversarial Low-Rank Adaptation of Vision-Language Models. arXiv:2404.13425.
- Jia et al. (2019) Xiaojun Jia, Xingxing Wei, et al. 2019. Comdefend: An efficient image compression model to defend adversarial examples. In CVPR.
- Jia et al. (2020) Xiaojun Jia, Xingxing Wei, et al. 2020. Adv-watermark: A novel watermark perturbation for adversarial examples. In ACM MM.
- Jiang et al. (2023) Albert Q Jiang, Alexandre Sablayrolles, Arthur Mensch, Chris Bamford, Devendra Singh Chaplot, et al. 2023. Mistral 7B. arXiv:2310.06825.
- Jigsaw (2017) Google Jigsaw. 2017. Perspective API. https://www.perspectiveapi.com/
- Jin et al. (2024) Haibo Jin et al. 2024. Jailbreakzoo: Survey, landscapes, and horizons in jailbreaking large language and vision-language models. arXiv:2407.01599.
- Jin et al. (2022) W Jin et al. 2022. A good prompt is worth millions of parameters: Low-resource prompt-based learning for vision-language models. In ACL.
@@
- Jordan et al. (2019) Matt Jordan, Naren Manoj, Surbhi Goel, et al. 2019. Quantifying perceptual distortion of adversarial examples. arXiv:1902.08265.
- Kannan et al. (2018) Harini Kannan, Alexey Kurakin, and Ian Goodfellow. 2018. Adversarial logit pairing. arXiv:1803.06373.
- Karli et al. (2022) Berat Tuna Karli, Deniz Sen, and Alptekin Temizel. 2022. Improving perceptual quality of adversarial images using perceptual distance minimization and normalized variance weighting. In AAAI Workshop.
- Kazemzadeh et al. (2014) Sahar Kazemzadeh, Vicente Ordonez, Mark Matten, et al. 2014. Referitgame: Referring to objects in photographs of natural scenes. In EMNLP.
@@
- Kingma (2013) Diederik P Kingma. 2013. Auto-encoding variational bayes. arXiv:1312.6114.
- Kingma (2014) Diederik P Kingma. 2014. Adam: A method for stochastic optimization. arXiv:1412.6980.
@@
- Krizhevsky et al. (2009) Alex Krizhevsky, Geoffrey Hinton, et al. 2009. Learning multiple layers of features from tiny images. Toronto, ON, Canada.
@@
- Kuo et al. (2025) Martin Kuo, Jianyi Zhang, Aolin Ding, et al. 2025. H-cot: Hijacking the chain-of-thought safety reasoning mechanism to jailbreak large reasoning models, including openai o1/o3, deepseek-r1, and gemini 2.0 flash thinking. arXiv:2502.12893.
- Kurakin et al. (2016) Alexey Kurakin, Ian Goodfellow, and Samy Bengio. 2016. Adversarial machine learning at scale. arXiv:1611.01236.
- Kurakin et al. (2018) Alexey Kurakin, Ian J Goodfellow, and Samy Bengio. 2018. Adversarial examples in the physical world. In Artificial intelligence safety and security.
- Laidlaw et al. (2020) Cassidy Laidlaw et al. 2020. Perceptual adversarial robustness: Defense against unseen threat models. arXiv:2006.12655.
@@
- LAION (2023) LAION. 2023. Reaching 80 zero-shot accuracy with openclip: Vit-g/14 trained on laion-2b. https://laion.ai/blog/giant-openclip/
- Lan et al. (2019) Xiangyuan Lan, Mang Ye, et al. 2019. Learning modality-consistency feature templates: A robust RGB-infrared tracking system. TIE.
- Larsen et al. (2016) Anders Boesen Lindbo Larsen, Søren Kaae Sønderby, et al. 2016. Autoencoding beyond pixels using a learned similarity metric. In ICML.
@@
- LeCun et al. (1998) Yann LeCun, Léon Bottou, Yoshua Bengio, and Patrick Haffner. 1998. Gradient-based learning applied to document recognition. Proc. IEEE.
- Lees et al. (2022) Alyssa Lees, Vinh Q Tran, Yi Tay, et al. 2022. A new generation of perspective api: Efficient multilingual character-level transformers. In SIGKDD.
- Li et al. (2023b) Bohao Li, Rui Wang, et al. 2023b. Seed-bench: Benchmarking multimodal llms with generative comprehension. arXiv:2307.16125.
- Li et al. (2023d) Bo Li, Yuanhan Zhang, Liangyu Chen, et al. 2023d. Otter: A Multi-Modal Model with In-Context Instruction Tuning. arXiv:2305.03726.
- Li et al. (2024d) Chao Li, Tingsong Jiang, et al. 2024d. Optimizing Latent Variables in Integrating Transfer and Query Based Attack Framework. TPAMI.
- Li et al. (2022b) Chao Li, Handing Wang, et al. 2022b. An approximated gradient sign method using differential evolution for black-box adversarial attack. TEC.
- Li et al. (2023c) Chao Li, Wen Yao, Handing Wang, Tingsong Jiang, and Xiaoya Zhang. 2023c. Bayesian evolutionary optimization for crafting high-quality adversarial examples with limited query budget. Applied Soft Computing.
- Li et al. (2018) Hao Li, Zheng Xu, Gavin Taylor, Christoph Studer, and Tom Goldstein. 2018. Visualizing the loss landscape of neural nets. In NIPS.
- Li et al. (2021) Junnan Li et al. 2021. Align before fuse: Vision and language representation learning with momentum distillation. In NIPS.
- Li et al. (2022a) Junnan Li et al. 2022a. Blip: Bootstrapping language-image pre-training for unified vision-language understanding and generation. In ICML.
- Li et al. (2023a) Junnan Li et al. 2023a. Blip-2: Bootstrapping language-image pre-training with frozen image encoders and large language models. In ICML.
- Li et al. (2019) Juncheng Li, Frank Schmidt, et al. 2019. Adversarial camera stickers: A physical camera-based attack on deep learning systems. In ICML.
- Li et al. (2024a) Lijun Li et al. 2024a. Salad-bench: A hierarchical and comprehensive safety benchmark for large language models. arXiv:2402.05044.
- Li et al. (2024c) Lin Li, Haoyan Guan, et al. 2024c. One prompt word is enough to boost adversarial robustness for pre-trained vision-language models. In CVPR.
- Li et al. (2020a) Maosen Li, Cheng Deng, Tengjiao Li, Junchi Yan, Xinbo Gao, and Heng Huang. 2020a. Towards transferable targeted attack. In CVPR.
- Li et al. (2024e) Mukai Li, Lei Li, Yuwei Yin, Masood Ahmed, Zhenguang Liu, and Qi Liu. 2024e. Red teaming visual language models. arXiv:2401.12915.
@@
- Li et al. (2020b) Yingwei Li et al. 2020b. Regional homogeneity: Towards learning transferable universal adversarial perturbations against defenses. In ECCV.
- Li et al. (2024b) Yifan Li et al. 2024b. Images are Achilles’ Heel of Alignment: Exploiting Visual Vulnerabilities for Jailbreaking Multimodal Large Language Models. arXiv:2403.09792.
- Li et al. (2025) Zhaoyi Li, Xiaohan Zhao, Dong-Dong Wu, Jiacheng Cui, and Zhiqiang Shen. 2025. A Frustratingly Simple Yet Highly Effective Attack Baseline: Over 90% Success Rate Against the Strong Black-box Models of GPT-4.5/4o/o1. arXiv:2503.10635.
- Liao et al. (2018) Fangzhou Liao, Ming Liang, et al. 2018. Defense against adversarial attacks using high-level representation guided denoiser. In CVPR.
@@
- Lin et al. (2020) Jiadong Lin, Chuanbiao Song, Kun He, et al. 2020. Nesterov accelerated gradient and scale invariance for adversarial attacks. ICLR.
- Lin et al. (2017) Yen-Chen Lin et al. 2017. Tactics of adversarial attack on deep reinforcement learning agents. arXiv:1703.06748.
- Lin et al. (2023) Zi Lin et al. 2023. Toxicchat: Unveiling hidden challenges of toxicity detection in real-world user-ai conversation. arXiv:2310.17389.
- Liu et al. (2020) Aishan Liu, Jiakai Wang, Xianglong Liu, Bowen Cao, et al. 2020. Bias-based universal adversarial patch attack for automatic check-out. In ECCV.
- Liu et al. (2024a) Daizong Liu et al. 2024a. A survey of attacks on large vision-language models: Resources, advances, and future trends. arXiv:2407.07403.
- Liu et al. (2023c) Fangcheng Liu, Chao Zhang, et al. 2023c. Towards transferable unrestricted adversarial examples with minimum changes. In SaTML.
- Liu et al. (2024b) Haotian Liu, Chunyuan Li, Yuheng Li, and Yong Jae Lee. 2024b. Improved baselines with visual instruction tuning. In CVPR.
- Liu et al. (2024c) Haotian Liu, Chunyuan Li, Yuheng Li, Bo Li, et al. 2024c. Llava-next: Improved reasoning, ocr, and world knowledge. https://llava-vl.github.io/blog/2024-01-30-llava-next/
- Liu et al. (2024d) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2024d. Visual instruction tuning. In NIPS.
- Liu et al. (2023d) X Liu, Y Zhu, J Gu, et al. 2023d. Mm-safetybench: A benchmark for safety evaluation of multimodal large language models. arXiv:2311.17600.
- Liu (2019) Yinhan Liu. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv:1907.11692.
- Liu et al. (2016) Yanpei Liu, Xinyun Chen, et al. 2016. Delving into transferable adversarial examples and black-box attacks. arXiv:1611.02770.
- Liu et al. (2022) Ye Liu, Yaya Cheng, Lianli Gao, Xianglong Liu, et al. 2022. Practical evaluation of adversarial robustness via adaptive auto attack. In CVPR.
- Liu et al. (2023a) Yi Liu, Gelei Deng, Zhengzi Xu, et al. 2023a. Jailbreaking chatgpt via prompt engineering: An empirical study. arXiv:2305.13860.
- Liu et al. (2023b) Yuan Liu, Haodong Duan, Yuanhan Zhang, et al. 2023b. Mmbench: Is your multi-modal model an all-around player? arXiv:2307.06281.
- Liu et al. (2021) Ze Liu, Yutong Lin, Yue Cao, Han Hu, et al. 2021. Swin transformer: Hierarchical vision transformer using shifted windows. In ICCV.
- Liu et al. (2019) Zihao Liu, Qi Liu, Tao Liu, Nuo Xu, et al. 2019. Feature distillation: Dnn-oriented jpeg compression against adversarial examples. In CVPR.
@@
- LMSYS (2023) LMSYS. 2023. Vicuna-7b-v1.5. https://huggingface.co/lmsys/vicuna-7b-v1.5
- Long et al. (2022) Yuyang Long, Qilong Zhang, Boheng Zeng, et al. 2022. Frequency domain model augmentation for adversarial attack. In ECCV.
- Lu et al. (2024) Dong Lu, Tianyu Pang, Chao Du, et al. 2024. Test-time backdoor attacks on multimodal large language models. arXiv:2402.08577.
- Luo et al. (2024a) H Luo et al. 2024a. An image is worth 1000 lies: Adversarial transferability across prompts on vision-language models. arXiv:2403.09766.
- Luo et al. (2024b) Weidi Luo et al. 2024b. Jailbreakv-28k: A benchmark for assessing the robustness of multimodal large language models against jailbreak attacks. arXiv:2404.03027.
- Ma et al. (2024) Siyuan Ma et al. 2024. Visual-RolePlay: Universal Jailbreak Attack on MultiModal Large Language Models via Role-playing Image Characte. arXiv:2405.20773.
- Madry (2017) Aleksander Madry. 2017. Towards deep learning models resistant to adversarial attacks. arXiv:1706.06083.
- Mao et al. (2016) Junhua Mao, Jonathan Huang, Alexander Toshev, et al. 2016. Generation and comprehension of unambiguous object descriptions. In CVPR.
@@
- Marino et al. (2019) Kenneth Marino, Mohammad Rastegari, et al. 2019. Ok-vqa: A visual question answering benchmark requiring external knowledge. In CVPR.
- Mazeika et al. (2024) Mantas Mazeika, Long Phan, et al. 2024. Harmbench: A standardized evaluation framework for automated red teaming and robust refusal. arXiv:2402.04249.
- Mazeika et al. (2023) Zou A. Mu N. Phan L. Mazeika, M. et al. 2023. Tdc 2023 (llm edition): The trojan detection challenge. In NIPS Competition Track.
- Microsoft (2023a) Microsoft. 2023a. Bing chat. https://www.bing.com/new
- Microsoft (2023b) Microsoft. 2023b. Copilot. https://copilot.microsoft.com/
- Mogelmose et al. (2012) Andreas Mogelmose, Mohan Manubhai Trivedi, and Thomas B Moeslund. 2012. Vision-based traffic sign detection and analysis for intelligent driver assistance systems: Perspectives and survey. TITS.
- Moosavi-Dezfooli et al. (2016) Seyed-Mohsen Moosavi-Dezfooli, Alhussein Fawzi, et al. 2016. Deepfool: a simple and accurate method to fool deep neural networks. In CVPR.
@@
- Mosaic (2023) Mosaic. 2023. Introducing MPT-7B: A New Standard for Open-Source, Commercially Usable LLMs. https://www.databricks.com/blog/mpt-7b
@@
- Narodytska and Kasiviswanathan (2016) Nina Narodytska and Shiva Prasad Kasiviswanathan. 2016. Simple black-box adversarial perturbations for deep networks. arXiv:1612.06299.
@@
- Naseer et al. (2020) Muzammal Naseer, Salman Khan, Munawar Hayat, et al. 2020. A self-supervised approach for adversarial robustness. In CVPR.
@@
- Nasr et al. (2023) Milad Nasr, Nicholas Carlini, et al. 2023. Scalable extraction of training data from (production) language models. arXiv:2311.17035.
@@
- Netzer et al. (2011) Yuval Netzer, Tao Wang, Adam Coates, et al. 2011. Reading digits in natural images with unsupervised feature learning. In NIPS Workshop.
@@
- Nguyen et al. (2015) Anh Nguyen et al. 2015. Deep neural networks are easily fooled: High confidence predictions for unrecognizable images. In CVPR.
@@
- Nie et al. (2022) Weili Nie, Brandon Guo, Yujia Huang, Chaowei Xiao, et al. 2022. Diffusion models for adversarial purification. arXiv:2205.07460.
- NIPS (2017) NIPS. 2017. NIPS 2017 Competition Track. https://nips.cc/Conferences/2017/CompetitionTrack